OpenStudy (anonymous):

Fundamental Theorem of Calculus: Find the derivative of h(x)=integral [1,e^x] ln(t) dt

myininaya (myininaya):

To integrate f(t) from t=1 to t=e^x we get F(x)=F(e^t)-F(1) Now to find the derivative of F we simply take the derivative of our above equation.

myininaya (myininaya):

You must know chain rule.

myininaya (myininaya):

that one t above is suppose to be an x

myininaya (myininaya):

F(x)=F(e^x)-F(1)

myininaya (myininaya):

The derivative of F(x) is of course (x)'*f(x)=1*f(x)=f(x) Now how do you find the derivative of F(e^x)?

OpenStudy (anonymous):

I have no idea what the integral of ln t is.

myininaya (myininaya):

ok I'm not asking you to do that

myininaya (myininaya):

I'm asking you to differentiate F(e^x) where F'=f

OpenStudy (anonymous):

isn't F supposed to be the antiderivative of the original function?

myininaya (myininaya):

I let f(t)=ln(t) I gave the antiderivative of f the name F(t) then I plugged in the upper limit minus plug in the lower limit giving me F(x)=F(e^x)-F(1) But we are suppose to be differentiating this so that is what I was asking you to differentiate F(e^x)

myininaya (myininaya):

\[\int\limits_{a}^{b}f(x) dx=F(x)|_a^b=F(b)-F(a) \text{ where F' =f } \] I'm using this here in this problem. \[\frac{d}{dx} (\int\limits_{1}^{e^x}f(t) dt )=\frac{d}{dx}(F(t)|_1^{e^x})=\frac{d}{dx}(F(e^x)-F(1))\]

myininaya (myininaya):

I'm replacing ln(t) with f(t) for right now we will come back and use it later

myininaya (myininaya):

I have to know. Do you understand this so far.

OpenStudy (anonymous):

Not really, but to be fair my brain is burned out. I'll have to come back to this, thanks for your time and explanations.

myininaya (myininaya):

I integrated so I got rid the integral sign

OpenStudy (anonymous):

I'm stuck on the formula and it just seems there is a creative approach you're using that's throwing me off. i'll come back to this thanks

myininaya (myininaya):

It isn't that creative.

myininaya (myininaya):

Do you have a problem with me calling the antiderivative of f F?

myininaya (myininaya):

Because that is all I really did so far